# Toward thermoelectric characterization of (nano)materials by in situ transmission electron microscopy

### v1

@@ Acknowledgements
The authors acknowledge funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 889546 and by the Spanish MCIN
(PID2019-104739GB-100/AEI/10.13039/501100011033 ) . The microscopy works have been conducted in the Laboratorio de Microscopias Avanzadas (LMA) at Universidad de Zaragoza. Sample courtesy (MLC NT) from MB Sreedhara and R. Tenne (MLC NT, Weizmann Institute of Science, Israel) is acknowledged. The authors thank M. Rengifo (INMA, Universidad de Zaragoza) for discussions about electrical characterizations, P. Strichovanek (INMA, Universidad de Zaragoza) for Ti-Pt sputter deposition and R. Valero (LMA, Universidad de Zaragoza) for support with MEMS fabrication.

### v2

@@ Acknowledgements
The authors acknowledge funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 889546, the Government of Aragon (DGA) through the project E13_23R, the Spanish MICIU with funding from European Union Next Generation EU (PRTR-C17.I1) promoted by the Government of Aragon and by the Spanish MICINN
(PID2019-104739GB-100/AEI/10.13039/501100011033 and CEX2023-001286-S MICIU/AEI /10.13039/501100011033 ) . The microscopy works have been conducted in the Laboratorio de Microscopias Avanzadas (LMA) at Universidad de Zaragoza. Sample courtesy (MLC NT) from MB Sreedhara and R. Tenne (MLC NT, Weizmann Institute of Science, Israel) is acknowledged. The authors thank M. Rengifo (INMA, CSIC-Universidad de Zaragoza) for discussions about electrical characterizations, P. Strichovanek (INMA, CSIC-Universidad de Zaragoza) for Ti-Pt sputter deposition, R. Valero (LMA, Universidad de Zaragoza) for support with MEMS and Ge thin-film fabrication and N. Navascues (INMA, CSIC-Universidad de Zaragoza) for support with Raman acquisition.